OpenStudy (electronia):

Brandon is bringing bottled water to a picnic. He has 4 full cases of bottles of water, and a bag with 8 more bottles. He has a total of 104 bottles. Write an equation to find the number of bottles in each case. Fan and metal for the first direct answer.

OpenStudy (kiakat):

okay so you want to find the amount of bottles in each case. You have four cases + 8 bottles right? All together that gives you 104 bottles. that would mean; cases + free bottles = total

OpenStudy (electronia):

Thats right! I having trouble mainly with what to do with 8. I will list the posible answers if that would be helpful.

OpenStudy (kiakat):

well 8 is just an addition. In order to find how many bottles are in the cases you'd have to subtract 8 from both sides of the equation cases + free bottles= total - free bottles - free bottles

OpenStudy (electronia):

Well, I need to amke an equation, so if I peice it together, would it be... x/4 + 8 =104 x=348?

OpenStudy (kiakat):

what are the possible answers?

OpenStudy (electronia):

Thats one, x/4 -8 = 104 x= 448 4x + 8= 104 x= 24 4x-8=104 x=28/

OpenStudy (kiakat):

okay. he has 4 cases PLUS 8 more bottles right?

OpenStudy (electronia):

Yes.

OpenStudy (kiakat):

that will automatically rule out the first and third option then. x/4 would mean the number of bottles in each case divided by four, which doesn't make much sense considering what you're trying to find. So it's definitely not the first option. 4x makes more sense because you have four cases, and x is the amount of bottles in each case. That would be 4 times the number of bottles in each case. +8 means he has 8 more bottles than what he has in the cases -8 means hes going to cut into the bottles that are already in the cases.

OpenStudy (electronia):

Thank you! Sorry for the time intake! Fanning!

OpenStudy (kiakat):

Not a problem! Happy to help ;v;